OpenStudy (anonymous):

Solid A is similar to Solid B with the given scale factor of A to B. The surface area and volume of Solid A are given. Find the surface area and volume of Solid B. Scale factor of 1:4 S = 62 cm squared V = 30 cm cubed

OpenStudy (anonymous):

www.lightandmatter.com/html_books/lm/ch01/ch01.html is the base of your question! ^^

OpenStudy (callisto):

\[(\frac{1}{4})^2 = \frac{62}{A}\]\[(\frac{1}{4})^3 = \frac{30}{V}\]

OpenStudy (anonymous):

after this do i just solve it?

Directrix (directrix):

Yes, for the first one: (1/4)^2 = 62/B as written above 1/16 = 62/B B = 16*62 = 992 square cm Work on the next equation while I find the statement of the theorem used to set up these equations.

Directrix (directrix):

(1/4)^3 = 30/V 1/64 = 30/V V = 64 * 30 V = 1920 cubic cm

OpenStudy (anonymous):

thank you Directrix :)